Find the intercepts and then use them to graph the equation. (I always have a hard time figuring out how to find the intercepts.)
y= -2-2x
Thank you

Answer by nerdybill(7384)

Given:
y= -2-2x
.
To find the x-intercepts, set y equal to zero and solve for x:
0= -2-2x
2 = -2x
-1 = x
so, x-intercept is at (-1,0)
.
To find the y-intercepts, set x equal to zero and solve for y:
y= -2-2(0
y= -2
so, y-intercept is at (0, -2)
.
To plot the line,
Plot each of the two points above and then simply draw a solid line through both points -- simple...